OpenStudy (anonymous):

*FAN AND MEDAL* help with Where are the asymptotes of f(x) = tan 4x from x = 0 to x =?

OpenStudy (thats_kyy):

Hold On Give Me A min Okay?

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

can you explain that ?? @Albany_Goon

OpenStudy (anonymous):

that all i got lemme do sum more digging

OpenStudy (anonymous):

okay

OpenStudy (freckles):

I don't understand why all of your questions end in to x=?

OpenStudy (anonymous):

i dont know either! theyre practice questions and im trying to review but ive never answered ones that end with x=?

OpenStudy (freckles):

anyways tan(4x)=sin(4x)/cos(4x) you just need to find when the bottom is 0 that is solve cos(4x)=0 for x

OpenStudy (freckles):

this will give you the vertical asymptotes of the graph of f(x)=tan(4x)

OpenStudy (freckles):

\[\cos(4x)=0 \\ \\ \text{ I guess we will just find them all } \\ \text{ first do you know how to solve } \\ \cos(u)=0\]

OpenStudy (anonymous):

no i dont

OpenStudy (freckles):

can you look on the unit circle and know for which angles the x coordinates are 0 http://managementscience.biz/wp-content/uploads/2014/10/unit-circle.gif

OpenStudy (freckles):

I will tell you one angle at 90 deg I see we have the x-coordinate is 0 so u=pi/2 rad is an answer to cos(u)=0 can you give another from looking at that picture?

OpenStudy (anonymous):

3pi/2?

OpenStudy (freckles):

awesome! so we have the solutions to: \[\cos(u)=0 \text{ are } u=\frac{\pi}{2}+2\pi n \text{ or } u=\frac{3\pi}{2}+2 \pi n \\ \text{ but we wanted to solve} \\ \cos(4x)=0 \\ \text{ well this equation is exactly the same as the one before except we had } \\ u \text{ instead of } 4x \\ \text{ just replace the } u's \text{ in the first equation } \\ \text{ with } 4x \\ \text{ that is you should have } \\ \cos(4x)=0 \text{ has solutions given by } \\ 4x=\frac{\pi}{2}+2 \pi n \text{ or } 4x=\frac{3\pi}{2}+2 \pi n \\ \] oh and n is an integer

OpenStudy (freckles):

you can finish solving for x by multiplying both sides by 1/4

OpenStudy (anonymous):

and then whatever i get after solving for x will be the asymptotes?

OpenStudy (freckles):

yep but make sure you leave the x in you know the asymptote should be written as an equation not just a value that is your answer should be x=something just take the equations I have given you and multiply both sides by 1/4

OpenStudy (freckles):

this will isolate the x

OpenStudy (anonymous):

can i do it and then you tell me if i get it right?

OpenStudy (freckles):

sure

OpenStudy (anonymous):

okay give me a minute please

OpenStudy (freckles):

ok I will be here as long as my internet doesn't mess up like it did earlier :p

OpenStudy (freckles):

been having some internet issues with my provider :(

OpenStudy (anonymous):

how do you multiply 1/4 times \[2\pi n\]

OpenStudy (anonymous):

and its fine i do to

OpenStudy (freckles):

\[4x=\frac{\pi}{2}+2 \pi n \text{ or } 4x=\frac{3\pi}{2}+2 \pi n \\ \text{ multiply both sides by } \frac{1}{4} \\ x=\frac{\pi}{2(4)}+\frac{2 \pi n}{4} \text{ or } x=\frac{3 \pi}{2(4)}+\frac{2 \pi n}{4} \\ \text{ simplifying } \\ x=\frac{\pi}{8}+\frac{\pi n}{2} \text{ or } x=\frac{3\pi}{8}+ \frac{\pi n }{2}\] note: since 2/4=1/2

OpenStudy (anonymous):

gotcha! thank you so much!!!

OpenStudy (freckles):

np